OpenStudy (anonymous):

Please help! A surveyor is “shooting a line” to a point on a tree 55 m from his current position. After rotating his surveying instrument 25° to the left, he “shoots” another line to a point on a fence post 25 m away.

OpenStudy (anonymous):

a.) Determine the distance between the point on the tree to the point on the fence post (TF). (Show the appropriate formula, substitutions, and work. Give the distance to the nearest hundredth of a meter.)

OpenStudy (anonymous):

b.) Use the Law of Sines to find the measure of T. (Show the appropriate formula, substitutions, and work. Give the measure of T to the nearest degree.)

OpenStudy (anonymous):

c.) Find the mF.

OpenStudy (anonymous):

@Thesmarterone @TheSmartOne

OpenStudy (alyssa_xo):

I think you can find FT using a trig relationship. sin=opp/hyp \(\sin(25)=\dfrac{x}{55}\)

OpenStudy (alyssa_xo):

so if that's right x=55*sin25 x=23.2

OpenStudy (triciaal):

@alyssa_xo are you sure SFT is 90 degrees?

OpenStudy (triciaal):

I think it would be obtuse not a right angle

OpenStudy (alyssa_xo):

oh, that was just an assumption, although given that it mentions the LOS and there isn't a square in that angle, I can't assume that

OpenStudy (triciaal):

start with the cosine rule c^2 = a^2 + b^2 -2abCos C

OpenStudy (triciaal):

then use the Sin rule sum of angles in a triangle = 180 or use the Sin rule again so now you have all lengths and all angles.

OpenStudy (triciaal):

to the nearest hundredth means to 2 decimal places calculate to 3 then round

OpenStudy (anonymous):

I got it figured out!

OpenStudy (triciaal):

great glad I could help